OpenStudy (anonymous):

when I find the deriative of (x-2)e^-x I am using the product rule and get -1. The answer is 1. I am missing a negative somewhere. Can you please help

OpenStudy (anonymous):

Show your work?

OpenStudy (anonymous):

(x-2)(-e^-x)+e^-x (1)

OpenStudy (anonymous):

mots likely you forgot to apply the chain rule on e^-x

OpenStudy (anonymous):

(d/dx)(e^-x) = -e^-x

OpenStudy (sidsiddhartha):

yeah \[\frac{ d }{ dx }e^{-ax}=-ae^{-ax}\]

OpenStudy (sidsiddhartha):

got that?

OpenStudy (anonymous):

got that but I am still getting -1. I applied the product rule and the chain rule for e^-x. Work [f'(x)= (x+2)(-e ^{-x})chain rule + e ^{-x}(1)] Product Rule [f'(x)=(-xe^-x)-(2e^-x)+(e^-x)] [f'(x)=(-xe^-x)-(e^-x)] I forgot to tell you I need to substitute 0 for x f'(0)=(-0e^-0)-(e^-0) f'(0)=(0) -(1) = -1 the answer is 1 please let me know how I missed a negative

OpenStudy (sidsiddhartha):

\[f(x)=(x-2)e^{-x}\] \[f'(x)=(x-2)[\frac{ d }{ dx}e^{-x}]+e^{-x}*\frac{ d }{ dx }(x-2)\] \[f'(x)=(x-2)[-e^{-x}]+e^{-x}=-e^{-x}[x-3]\] ok so far?

OpenStudy (sidsiddhartha):

now if u put x=0 then \[f'(0)=-e^{0}[0-3]=3\] thats ur answer it is not (1)

OpenStudy (anonymous):

Thank you. You showed me how my Algebra was off. The problem was f(x)=(x+2)e^-x x=0. You provided me f(x)=(x-2)e^-x but I was still easily able to follow and got my answer of 1. Thanks again